OpenStudy (anonymous):

Estimate by first rounding each number to the nearest 1/2 -3 8/9 * 1 1/2 A. –6 B. –3 C. 6 D. 8

OpenStudy (anonymous):

put ur question again

OpenStudy (anonymous):

Estimate by first rounding each number to the nearest \[\frac{ 1 }{ 2} \] \[-3\left(\begin{matrix}8 \\ 9\end{matrix}\right)\] \[\times 1 \left(\begin{matrix}1 \\ 2\end{matrix}\right)\] A. –6 B. –3 C. 6 D. 8